# DualPM: Dual Posed-Canonical Point Maps for 3D Shape and Pose Reconstruction

### v3

@@ Abstract
The choice of data representation is a key factor in the success of deep learning in geometric tasks. For instance, DUSt3R has recently introduced the concept of viewpoint-invariant point maps, generalizing depth prediction, and showing that one can reduce all the key problems in the 3D reconstruction of static scenes to predicting such point maps. In this paper, we develop an analogous concept for a very different problem, namely, the reconstruction of the 3D shape and pose of deformable objects. To this end, we introduce the Dual Point Maps (DualPM), where a pair of point maps is extracted from the same image, one associating pixels to their 3D locations on the object, and the other to a canonical version of the object at rest pose. We also extend point maps to amodal reconstruction, seeing through self-occlusions to obtain the complete shape of the object. We show that 3D reconstruction and 3D pose estimation reduce to the prediction of the DualPMs. We demonstrate empirically that this representation is a good target for a deep network to predict; specifically, we consider modeling horses, showing that DualPMs can be trained purely on 3D synthetic data, consisting of a single model of a horse, while generalizing very well to real images. With this, we improve by a large margin previous methods for the 3D analysis and reconstruction of this type of objects.
@@ 1 Introduction
Until recently, the general consensus was that inferring accurate 3D information should be, whenever possible, left to visual geometry. Hence, neural networks would fill gaps such as feature matching and monocular prediction where visual geometry is not applicable. However, this view is increasingly challenged as authors find better ways of reducing geometric tasks to the calculation of quantities that can be predicted very effectively by neural networks. Most recently, DUSt3R [65, 8] has demonstrated the power of predicting point maps, namely images that associate to each pixel its corresponding 3D point in the scene. They show that many tasks in the reconstruction of static scenes, such as matching, camera estimation, and triangulation, can be solved trivially from the point maps predicted by a neural network.
In this work, we ask whether a similar intuition is applicable to a very different class of problems in visual geometry, namely the monocular reconstruction of the 3D shape and pose of deformable objects. We do so by developing a new network-friendly representation, which we call Dual Point Maps (DualPMs).
Our setting is a substantial departure from DUSt3R, and so are their point maps and our DualPMs. Nevertheless, the starting point is the same: a point map in both DUSt3R and DualPM associates to each image pixel its corresponding 3D point on the object. Predicting this point map gives us a 3D reconstruction of the object, which is useful, but gives us no information about the pose of the object.
To clarify this further, consider the example of a horse. Its pose is defined as the deformation that takes the horse from its neutral pose to the pose observed in the image. Hence, reconstructing its pose means finding the transformation, or deformation field, between two versions of the 3D object: the one seen in the image and the canonical version of the same, in neutral pose. Knowledge of a single point map gives no information about this transformation. So the question is: How can we extend point maps so that recovering this deformation field becomes trivial?
To solve this problem, our DualPMs predicts not one, but two point maps from the same image (Fig. 1). They map each pixel to two versions of the same 3D point. The first version is the usual reconstruction of the object in 3D. The second version, however, is the location of the point in the canonical version of the object. With this representation, it is easy to recover the pose of the object: in fact, the deformation field that we wish to recover is simply the difference between the two point maps. Knowledge of this deformation field is in turn useful for various analysis tasks, such as recovering 3D keypoints or estimating the object’s articulation in terms of a skeleton (see  Fig. 1).
@@
A shortcoming of point maps, in general, is that they only reconstruct the visible part of the object. To address this issue, we further introduce the idea of estimating amodal point maps; namely, each pixel is mapped not only to the visible point of the object but to all points on the object that project onto that pixel, regardless of self-occlusions (Fig. 2). To effectively represent such point maps for neural network prediction, we propose a layered representation. In this representation, the first layer encodes the visible points, and each subsequent layer captures the next set of points occluded by the previous layer. This concept is similar to depth peeling, a standard technique used in rasterization in computer graphics. With this extension, we can provide a full reconstruction of the object’s shape and pose.
To test our ideas, we consider the problem of reconstructing the 3D shape and pose of horses from monocular images. We do so as horses represent a widely studied category of deformable objects [25, 69, 71, 17, 32], with synthetic 3D data readily available for this class [17], enabling controlled evaluations. Furthermore, we show that synthetic data derived from a single horse model is sufficient to train a DualPM model hat generalizes robustly to real-world data thanks to our effective representation that makes learning relatively easy. With this, we show that we can improve previous methods by a large margin for the 3D analysis and reconstruction of this type of objects, even on of out-of-domain real-world datasets, both in terms of correspondence and 3D reconstruction. Moreover, our method exhibits strong zero-shot generalization to real-world images of other quadruped animals, such as cow and sheep, despite being trained exclusively on a single horse model. We also demonstrate the ability of the point maps to reconstruct the pose of the object in terms of their deformation fields, and suggest applications such as fitting articulated 3D skeletons to the object that can be used for motion transfer and animations.
@@
    We introduce the novel concept of DualPM, which reduces monocular 3D shape and pose reconstruction to the prediction of a pair of point maps from the image: one defined in canonical space and the other in the posed camera space.
@@
    We extend point maps to enable amodal reconstruction along with their efficient layered representation, producing the complete 3D shape of the articulated object.
@@
    We demonstrate that DualPMs can be effectively predicted by a neural network and that this representation serves as a suitable target for learning, to the extent that synthetic data is sufficient for training the model.
@@
    We show key tasks in the analysis of deformable objects, including 3D reconstruction, keypoint transfer/matching, estimation of deformation, and fitting of articulated skeletons, reduce to the prediction of DualPMs.
@@
    We demonstrate that DualPMs significantly outperform prior approaches for 3D analysis and reconstruction of horses and generalize effectively to other quadrupeds, even when trained exclusively on synthetic data from a single articulated horse model.
@@ 2 Related work
Our work is related to the reconstruction of deforming scenes. Many such works consider the optimization approach, where a 4D model of the scene is fitted to a video sequence. In works like Video Pop-up[50], Monocular 3D [26], MonoRec [66] the 3D reconstruction of different video frames are regularized, for instance to enforce approximate local rigidity.
More often, the scene 4D scene is represented as the deformation of a canonical configuration. DynamicFusion [42] is an early example of this approach. More recent works like Neural Volumes [39] reconstruct both shape and appearance. D-NeRF [46], Neural Radiance Flow [7], Neural Scene Flow Fields [30], Dynamic Video Synthesis [15], DynIBaR [31], MorpheuS [60], extend NeRF [41] with dense deformation fields to capture motion. Similarly, 4D Gaussian Splatting [67], Gaussian Flow [35], DynMF [23], GauFRe [33], MoSca [27], Street Gaussians [73], Dynamic Gaussian Marbles [55], Shape of Motion [64], Ego Gaussians [83], 4DGS [67] also estimate deformations, but in the context of 3D Gaussian Splatting [20] rather than NeRF.
Rather than modelling dense deformation fields, approaches like Dynamic NeRF [14], Seeing 3D Objects [51], K-planes [13] and HexPlane [4] just add time or pose parameterization to the NeRF model. Approaches like 4D Visualization [2] and NVSD [82] predict novel views of a dynamic scene without explicit 3D reconstruction.
Fourier PlenOctrees [63], NeRFPlayer [53] and RT-4DGS [78] focus on the problem of storing and rendering 4D models efficiently. 4DGen [81], Align Your Gaussians [36], DreamScene4D [6], L4GM [47] considered instead the problem of generating 4D models from text (text-to-4D). PhysGaussian [72], MD-Splatting [9], Gaussian Splashing [12] and VR-GS [18] consider reconstruction and generation of 4D scene which accounts for physical principles
Works like Neural Human Video Rendering [38], AutoAvatar [1], Neural Body [45], Dynamic Facial RF [52], Relighting4D [5], Animate124 [86], Dynamic Gaussian Mesh [37] and IM4D [34] specialise in the reconstruction of articulated characters like humans and animals. Often these define the object in a canonical neutral pose, and encode pose as a deformation applied to this pose.
@@
Works like 3D Menagerie [87], Unsupervised 3D [68], LASR [74], BANMo [75], Dessie [28], DOVE [69], LASSIE [79], MagicPony [70], Farm3D [17], 3D Fauna [32], and others [16, 88, 76, 80, 49] have considered the problem of learning models of 3D object categories. In these models, pose is expressed as the object deformation with respect to a version of the object in neutral pose, which usually takes the form of a category-level template. Such a template can be learned either in a weakly-supervised manner [69, 79, 70, 17, 32, 76, 80] or from 3D data, most notably SMPL [40], SMAL [3], and more recently, VAREN [89]. Concurrent Dessie also introduces a synthetic data generation pipeline similar to ours.
@@
DUSt3R [65] was the first paper to intuit the power of predicting point maps instead of depth for 3D reconstruction. The follow up MASt3R [8] builds on this to obtain a state-of-the-art Structure-from-Motion (SfM) system. MonST3R [85] extend DUSt3R to monocular video, reconstructing dynamic point maps in a fixed reference frame.
The concept of learning a map that sends pixels to canonical object points can be found in the dense equivariant mapping of [56, 57, 58]. Normalized object coordinates NOCs [59] modify this idea for rigid objects by sending pixels to 3D points defined in the object’s surface in canonical pose. Our approach also does this, but with two key differences. First, the input object is deformable rather than rigid. Second, we predict dual point maps, which encode the object deformation beyond a simple rigid transformation.
@@ 3 Method
Let $I\in\mathbb{R}^{3\times H\times W}$ be an image containing a deformable object we wish to reconstruct in 3D, along with its pose. We assume that a mask $M\in\{0,1\}^{H\times W}$ is available, indicating the object’s location in the image. Next, we introduce the concept of dual point maps (Sec. 3.1) and then of its amodal extension (Sec. 3.2).
@@ 3.1 Dual point maps
A point map $P\in\mathbb{R}^{3\times H\times W}$ is a tensor containing a 3D point for each pixel of the input image $I$. We are, in particular, interested in the pixels $\boldsymbol{u}\in M$ contained within the object’s mask $M$. Then $\boldsymbol{u}$ is the image of some point $\boldsymbol{p}$ that belongs to the surface of the object, known as the imaged point. The image sample $I(\boldsymbol{u})$ is the color of the imaged point (an RGB triplet), and $\boldsymbol{p}=P(\boldsymbol{u})\in\mathbb{R}^{3}$ is the 3D location of the point, which we express in the reference frame of the camera.
By estimating the point map $P$ from the image $I$, we reconstruct the visible portion of the 3D object. This is similar to predicting a depth map but provides strictly more information. In fact, the 3D point does not depend on the camera’s intrinsic parameters, such as the focal length, as it is directly expressed in the camera’s reference frame. As shown in [65], this fact can be used to infer the camera intrinsics. However, this point map does not capture the object’s deformation and thus cannot be used by itself to recover its pose.
We address this limitation by introducing the novel concept of Dual Point Maps (DualPMs). In addition to $P$, we predict a second point map $Q\in\mathbb{R}^{3\times H\times W}$ defined in canonical space. By canonical, we mean two things: first, the points in $Q$ are expressed in the object reference frame rather than in the camera one; second, the object is in a canonical pose, usually a neutral pose. Hence, differently from $P$, the point map $Q$ is invariant to the object’s pose and deformation.
The pair of point maps $P$ and $Q$ implicitly encode the object deformation: they tell us that pixel $\boldsymbol{u}$ corresponds to a point which is found at location $\boldsymbol{q}=Q(\boldsymbol{u})$ in rest pose and at location $\boldsymbol{p}=P(\boldsymbol{u})$ once posed. This provides rich information about the object pose. For example, given two images $I_{1}$ and $I_{2}$ of the object in two different poses, we can match a pixel $\boldsymbol{u}_{1}$ in the first image to its corresponding pixel $\boldsymbol{u}_{2}$ in the second by matching the corresponding canonical points $Q_{1}(\boldsymbol{u}_{1})$ and $Q(\boldsymbol{u}_{2})$, i.e.,
@@
This works because the canonical points are view and pose invariant. Then, we can infer the so-called ‘scene flow’, or motion between the 3D points, as $P_{2}(\boldsymbol{u}_{2}(\boldsymbol{u}_{1}))-P_{1}(\boldsymbol{u}_{1}).$
@@
To predict the dual point maps from the image $I$, we introduce a predictor neural network $(P,Q)=\Phi(I).$ Its architecture is illustrated in Fig. 3. A key property of the canonical point map $Q$ is its invariance to the object’s pose and deformation. This invariance simplifies its prediction by a neural network, as it reduces it to a pixel labeling problem. To further ease the prediction of $Q$ and improve generalization, we leverage recent advances in self-supervised image feature learning [84, 44, 10], and condition its prediction on strong features $F$ extracted from the image $I$ using a pre-trained network $\Psi$, $F=\Psi(I)$ from [84]. These features serve as a good proxy for the canonical point map, as they have been shown to be nearly invariant to the object’s pose and deformation, thereby simplifying the network’s task. We then condition the prediction of $P$ on $Q$ instead of $F$ as it improves generalization of the model on out-of-distribution images, as demonstrated in Sec. 4.5. This shows that learning canonical point maps as an intermediate representation benefits the 3D reconstruction task. To summarize, we first predict $Q$ using the features $F$, and then condition the prediction of $P$ on $Q$:
@@
We extend both networks $\Phi_{Q}$ and $\Phi_{P}$ to also predict a per-pixel confidence score. So, for example, the map $\Phi_{P}$ outputs the map $c_{P}(\boldsymbol{u})>0$, which is used to train the model by means of the self-calibrated L2 loss of [43, 19, 65]:
@@
where $\alpha>0$ is a constant and $\hat{P}$ is the ground-truth point map. The analogous loss $\mathcal{L}_{Q}$ is used for $\Phi_{Q}$.

### v4

@@ Abstract
The choice of data representation is a key factor in the success of deep learning in geometric tasks. For instance, DUSt3R recently introduced the concept of viewpoint-invariant point maps, generalizing depth prediction and showing that all key problems in the 3D reconstruction of static scenes can be reduced to predicting such point maps. In this paper, we develop an analogous concept for a very different problem: the reconstruction of the 3D shape and pose of deformable objects. To this end, we introduce Dual Point Maps (DualPM), where a pair of point maps is extracted from the same image—one associating pixels to their 3D locations on the object and the other to a canonical version of the object in its rest pose. We also extend point maps to amodal reconstruction to recover the complete shape of the object, even through self-occlusions. We show that 3D reconstruction and 3D pose estimation can be reduced to the prediction of DualPMs. Empirically, we demonstrate that this representation is a suitable target for deep networks to predict. Specifically, we focus on modeling quadrupeds, showing that DualPMs can be trained purely on synthetic 3D data, consisting of one or two models per category, while generalizing effectively to real images. With this approach, we achieve significant improvements over previous methods for the 3D analysis and reconstruction of such objects.
@@ 1 Introduction
Until recently, the general consensus was that inferring accurate 3D information should be, whenever possible, left to visual geometry, leaving to neural networks to fill gaps such as feature matching and monocular prediction. However, this view is increasingly challenged as researchers find better ways of reducing geometric tasks to the calculation of quantities that can be predicted very effectively by neural networks. Most recently, DUSt3R [66, 9] has demonstrated the power of predicting point maps, namely images that associate each pixel with its corresponding 3D point in the scene. They show that many tasks in the reconstruction of static scenes, such as matching, camera estimation, and triangulation, can be solved trivially from the point maps predicted by a neural network.
In this work, we ask whether a similar intuition applies to a different class of problems, namely the monocular reconstruction of the 3D shape and pose of deformable objects. We do so by developing a new network-friendly representation, which we call Dual Point Maps (DualPMs).
Our setting is a substantial departure from DUSt3R, and so are their point maps and our DualPMs. Nevertheless, the starting point is the same: a point map in both DUSt3R and DualPM associates each image pixel with its corresponding 3D point on the object. Predicting this point map gives us a 3D reconstruction of the object, which is useful but provides no information about the pose of the object.
To clarify this further, consider the example of a horse. Its pose is defined as the deformation that takes the horse from its neutral pose to the pose observed in the image. Hence, reconstructing its pose means finding the transformation, or deformation field, between two versions of the 3D object: the one seen in the image and the canonical version of the same object in a neutral pose. Knowledge of a single point map gives no information about this transformation. So the question is: How can we extend point maps so that recovering this deformation field becomes trivial?
To solve this problem, our DualPMs predicts not one, but two point maps from the same image (Fig. 1). They map each pixel to two versions of the same 3D point. The first version is the usual reconstruction of the object in 3D. The second version, however, is the location of the point in the canonical version of the object. With this representation, it is easy to recover the pose of the object: in fact, the deformation field that we wish to recover is simply the difference between the two point maps. Knowledge of this deformation field is, in turn, useful for various analysis tasks, such as recovering 3D keypoints or estimating the object’s articulation in terms of a skeleton (see Fig. 1).
@@
A shortcoming of point maps, in general, is that they only reconstruct the visible part of the object. To address this issue, we introduce the idea of estimating amodal point maps; namely, each pixel is mapped not only to the visible point of the object but to all points on the object that project onto that pixel, regardless of self-occlusions (Fig. 2). To effectively represent such point maps for neural network prediction, we propose a layered representation. In this representation, the first layer encodes the visible points, and each subsequent layer captures the next set of points occluded by the previous layer. This concept is similar to depth peeling, a standard technique used in rasterization in computer graphics. With this extension, we can provide a full reconstruction of the object’s shape and pose.
To test our ideas, we consider the problem of reconstructing the 3D shape and pose of quadrupeds from monocular images. Quadrupeds represent a widely studied category of deformable objects [26, 71, 72, 18, 33], with synthetic 3D data readily available for this class [18], enabling controlled evaluations. Furthermore, we show that synthetic data derived from one or two 3D models per category is sufficient to train a DualPM model that generalizes robustly to real-world data thanks to our effective representation, which makes learning relatively easy. With this, we demonstrate that we can improve previous methods by a large margin for the 3D analysis and reconstruction of this type of object, even on out-of-domain real-world datasets, both in terms of correspondence and 3D reconstruction. We also show the ability of the point maps to reconstruct the pose of the object in terms of their deformation fields and suggest applications such as fitting articulated 3D skeletons to the object, which can be used for motion transfer and animations.
@@
    We introduce the novel concept of DualPM, reducing monocular 3D shape and pose reconstruction to predicting a pair of point maps: one in canonical space and the other in posed camera space.
@@
    We extend point maps to enable amodal reconstruction with a layered representation, producing the complete 3D shape of articulated objects.
@@
    We demonstrate that DualPMs can be effectively predicted by neural networks, with synthetic data sufficing for training.
@@
    We show that tasks like 3D reconstruction, keypoint transfer, deformation estimation, and skeleton fitting reduce to predicting DualPMs.
@@
    We achieve significant improvements over prior methods for 3D analysis and reconstruction of quadrupeds, even when trained on minimal synthetic data.
@@ 2 Related work
Our work is related to the reconstruction of deforming scenes. Many such works consider the optimization approach, where a 4D model of the scene is fitted to a video sequence. In works like Video Pop-up [51], Monocular 3D [27], and MonoRec [67], the 3D reconstruction of different video frames is regularized, for instance, to enforce approximate local rigidity.
More often, the 4D scene is represented as the deformation of a canonical configuration. DynamicFusion [43] is an early example of this approach. More recent works like Neural Volumes [40] reconstruct both shape and appearance. D-NeRF [47], Neural Radiance Flow [8], Neural Scene Flow Fields [31], Dynamic Video Synthesis [16], DynIBaR [32], and MorpheuS [61] extend NeRF [42] with dense deformation fields to capture motion. Similarly, 4D Gaussian Splatting [68], Gaussian Flow [36], DynMF [24], GauFRe [34], MoSca [28], Street Gaussians [75], Dynamic Gaussian Marbles [56], Shape of Motion [65], Ego Gaussians [85], and 4DGS [68] also estimate deformations, but in the context of 3D Gaussian Splatting [21] rather than NeRF.
Rather than modeling dense deformation fields, approaches like Dynamic NeRF [15], Seeing 3D Objects [52], K-planes [14], and HexPlane [5] add time or pose parameterization to the NeRF model. Approaches like 4D Visualization [2] and NVSD [84] predict novel views of a dynamic scene without explicit 3D reconstruction.
Fourier PlenOctrees [64], NeRFPlayer [54], and RT-4DGS [80] focus on the problem of storing and rendering 4D models efficiently. 4DGen [83], Align Your Gaussians [37], DreamScene4D [7], and L4GM [48] consider the problem of generating 4D models from text (text-to-4D). PhysGaussian [74], MD-Splatting [10], Gaussian Splashing [13], and VR-GS [19] consider reconstruction and generation of 4D scenes that account for physical principles.
Works like Neural Human Video Rendering [39], AutoAvatar [1], Neural Body [46], Dynamic Facial RF [53], Relighting4D [6], Animate124 [88], Dynamic Gaussian Mesh [38], and IM4D [35] specialize in the reconstruction of articulated characters like humans and animals. Often, these define the object in a canonical neutral pose and encode the pose as a deformation thereof.
@@
Works like 3D Menagerie [89], Unsupervised 3D [69], LASR [76], BANMo [77], Dessie [29], DOVE [71], LASSIE [81], MagicPony [70], Farm3D [18], 3D-Fauna [33], and others [17, 90, 78, 82, 50] have considered the problem of learning models of 3D object categories. In these models, the pose is expressed as the object deformation with respect to a version of the object in a neutral pose, which usually takes the form of a category-level template. Such a template can be learned either in a weakly-supervised manner [71, 81, 70, 18, 33, 78, 82] or from 3D data, most notably SMPL [41], SMAL [3], and more recently, VAREN [91]. Concurrent Dessie also introduces a synthetic data generation pipeline similar to ours.
@@
DUSt3R [66] was the first paper to intuit the power of predicting point maps instead of depth for 3D reconstruction. The follow-up MASt3R [9] builds on this to obtain a state-of-the-art Structure-from-Motion (SfM) system. MonST3R [87] extends DUSt3R to monocular video, reconstructing dynamic point maps in a fixed reference frame.
The concept of learning a map that sends pixels to canonical object points can be found in the dense equivariant mapping of [57, 58, 59]. Normalized Object Coordinates (NOCs) [60] modify this idea for rigid objects by sending pixels to 3D points defined on the object’s surface in a canonical pose. Our approach also does this but with two key differences: (1) the input object is deformable rather than rigid, and (2) we predict dual point maps, which encode the object deformation beyond a simple rigid transformation.
@@ 3 Method
Let $I\in\mathbb{R}^{3\times H\times W}$ be an image containing a deformable object we wish to reconstruct in 3D, along with its pose. We assume that a mask $M\in\{0,1\}^{H\times W}$ is available, indicating the object’s location in the image. Next, we introduce the concept of dual point maps (Sec. 3.1) and its amodal extension (Sec. 3.2).
@@ 3.1 Dual point maps
A point map $P\in\mathbb{R}^{3\times H\times W}$ is a tensor containing a 3D point for each pixel of the input image $I$. We are particularly interested in the pixels $\boldsymbol{u}\in M$ contained within the object’s mask $M$. Here, $\boldsymbol{u}$ is the image of some point $\boldsymbol{p}$ that belongs to the surface of the object, known as the imaged point. The image sample $I(\boldsymbol{u})$ is the color of the imaged point (an RGB triplet), and $\boldsymbol{p}=P(\boldsymbol{u})\in\mathbb{R}^{3}$ is the 3D location of the point, expressed in the reference frame of the camera.
By estimating the point map $P$ from the image $I$, we reconstruct the visible portion of the 3D object. This is similar to predicting a depth map but provides strictly more information. In fact, the 3D point does not depend on the camera’s intrinsic parameters, such as the focal length, as it is directly expressed in the camera’s reference frame. As shown in [66], this fact can be used to infer the camera intrinsics. However, this point map does not capture the object’s deformation and thus cannot be used by itself to recover its articulated pose.
We address this limitation by introducing the novel concept of Dual Point Maps (DualPMs). In addition to $P$, we predict a second point map $Q\in\mathbb{R}^{3\times H\times W}$ defined in canonical space. By “canonical”, we mean two things: first, the points in $Q$ are expressed in the object reference frame rather than in the camera frame; second, the object is in a canonical pose, usually a neutral pose. Hence, unlike $P$, the point map $Q$ is invariant to the object’s pose and deformation.
The pair of point maps $P$ and $Q$ implicitly encode the object deformation: they tell us that pixel $\boldsymbol{u}$ corresponds to a point that is located at $\boldsymbol{q}=Q(\boldsymbol{u})$ in the rest pose and at $\boldsymbol{p}=P(\boldsymbol{u})$ once posed. This provides rich information about the object’s pose. For example, given two images $I_{1}$ and $I_{2}$ of the object in two different poses, we can match a pixel $\boldsymbol{u}_{1}$ in the first image to its corresponding pixel $\boldsymbol{u}_{2}$ in the second by matching the corresponding canonical points $Q_{1}(\boldsymbol{u}_{1})$ and $Q_{2}(\boldsymbol{u}_{2})$, i.e.,
@@
This works because the canonical points are view- and pose-invariant. Then, we can infer the so-called ‘scene flow’, or motion between the 3D points, as $P_{2}(\boldsymbol{u}_{2}(\boldsymbol{u}_{1}))-P_{1}(\boldsymbol{u}_{1}).$
@@
To predict the dual point maps from the image $I$, we introduce a predictor neural network $(P,Q)=\Phi(I).$ Its architecture is illustrated in Fig. 3. A key property of the canonical point map $Q$ is its invariance to the object’s pose and deformation. This invariance simplifies its prediction by a neural network, as it reduces the task to a pixel labeling problem. To further ease the prediction of $Q$ and improve generalization, we leverage recent advances in self-supervised image feature learning [86, 45, 11], and condition its prediction on strong features $F$ extracted from the image $I$ using a pre-trained network $\Psi$, $F=\Psi(I)$ from [86]. These features serve as a good proxy for the canonical point map, as they have been shown to be nearly invariant to the object’s pose and deformation, thereby simplifying the network’s task. We then condition the prediction of $P$ on $Q$ instead of $F$, as this improves the model’s generalization on out-of-distribution images, as demonstrated in Sec. 4.4. This shows that learning canonical point maps as an intermediate representation benefits the 3D reconstruction task. To summarize, we first predict $Q$ using the features $F$, and then condition the prediction of $P$ on $Q$:
@@
We extend both networks $\Phi_{Q}$ and $\Phi_{P}$ to also predict a per-pixel confidence score. For example, the map $\Phi_{P}$ outputs the map $c_{P}(\boldsymbol{u})>0$, which is used to train the model using the self-calibrated L2 loss from [44, 20, 66]:
@@
where $\alpha>0$ is a constant and $\hat{P}$ is the ground-truth point map. An analogous loss $\mathcal{L}_{Q}$ is used for $\Phi_{Q}$.